OpenStudy (briggles1999):

which equation represents the line that passes through the points (6 –3) and (–4 –9) y+4=3/5(x+9) y+4=5/3(x+9) y-3=3/5(x+6) y+3=3/5(x-6)

OpenStudy (shinalcantara):

To get the equation of the line, find first the slope of the line \[m = \frac{y_2-y_1}{x_2-x_1}\] \[m = \frac{ -9-(-3) }{ -4-6 }\]

OpenStudy (briggles1999):

??????

OpenStudy (briggles1999):

@shinalcantara

OpenStudy (shinalcantara):

\[m = \frac{ 3 }{ 5 }\] Using point-slope form and one of the points say (6,-3) (y-y1)=m(x-x1) (y-(-3))=3/5(x-6) y+3 = 3/5(x-6)